### v1

@@ II-A4 Usage Snippets
To help the LLM generate a valid exploit, PoCGen extracts usage examples of the vulnerable function from the codebase. These examples allow the LLM to understand the function signature, and any setup that is required to call the function. We extract usage snippets both from the source code and from the documentation of the package. The usage snippets from the source code are extracted from test files using static analysis, by finding all the call sites of the vulnerable function. For the usage snippets from the documentation, we first extract all code pieces in the documentation, wrapped in triple backticks (‘‘‘), and then we prompt an LLM to determine if they are usage examples of the vulnerable function. If they are, we also prompt the LLM to summarize them.
II-B Exploit Generation
@@
PoCGen assembles a prompt for the LLM to generate the PoC exploit. Figure 5 shows the prompt template that we use for generating the exploit. The prompt starts with naming the vulnerable function (as vulnerableFunction) and the vulnerability type (as vulnerabilityType), which are extracted in the previous phase, followed by a description of the vulnerability (as vulnerabilityDescription), which is the vulnerability report text. Then, we provide some example usages of the vulnerable function (as usageSnippets), which are extracted from the codebase. In the last section of the prompt, we describe how the generated exploit should look like and what it should do, by providing the skeleton of the exploit code (as exploitSkeleton), and also providing exploits of similar vulnerabilities (as similarExploits). The skeleton of the exploit code, is a fixed template that contains the definition of the exploit function, its call, and loading the vulnerable function from the vulnerable package. To extract similar exploits, our approach uses BM25 to find the three most similar vulnerability descriptions from the SecBench.js dataset to the current vulnerability report, and provides their exploits as examples.
Figure 5: Prompt template for the first round of exploit generation.
{minted}
[breaksymbol=]handlebars ‘vulnerableFunction‘ is vulnerable to vulnerabilityType. ## Vulnerability Description: “‘ vulnerabilityDescription “‘ ¿ usageSnippets ¿ exploitSkeleton ## Task: Create an exploit that uses the vulnerabilityType in ‘vulnerableFunction‘ to vulnerabilityTrigger. Respond with the full exploit code and explain why it works. If there is an exception thrown, do not try to handle it and pass it on. Enclose the exploit code in backticks and define the exploit within a function named ‘exploit‘. ¿ similarExploits ## Source code: taintPathSnippets
II-C Exploit Validation
@@
We define the goal of the exploit specific to each vulnerability type. This allows us to validate the generated exploit with simple checkers. These checkers are inspired by the validation checks in SecBench.js [5].
Path Traversal.
@@
Prototype Pollution.
In the case of prototype pollution vulnerabilities, the goal of the exploit is to add a property named exploited to the global Object.prototype object. However, this does not always guarantee that the exploit is valid, as the property can be added to the object without exploiting the vulnerability. To filter out such cases, we also check if the expression of the form *[__proto__|prototype].exploited = * is present in the exploit code. If such pattern exists, the exploit is considered invalid.
Command Injection.
@@
II-D Prompt Refinement
After the validation step, if the exploit is not valid and the maximum number of refinements is not exceeded, PoCGen refines the prompt to generate a new candidate exploit. PoCGen uses a set of refiners that provide static or dynamic information to the LLM to help it generate a valid exploit.
Context Refiners.
@@ II-D Prompt Refinement
However, there can be checks that happen via function calls that are not in the taint path. To address this, we also provide a missing declaration refiner, which provides the LLM with the ability to ask for definitions of variables and functions in the taint path, through the function calling format of OpenAI’s API. The LLM can output a list of identifiers that it needs their definitions, and PoCGen will provide the definitions of these identifiers in the prompt.
Runtime Refiners.
@@
The debugger refiner provides the LLM with a debugger-like tool. The LLM can output a list of expressions, for which it needs the runtime values. The refiner will provide the values of these expressions during the execution of the exploit in the prompt. These values are provided as comments in their respective lines in the taint path.
@@
In every refinement attempt, PoCGen chooses one refiner from the front of a priority queue. Initially all refiners are in the queue. Each time the approach uses a refiner, it assigns a score based on the number of new errors the respective exploit causes, and the number of steps from the taint path it covers. It then adds the refiner to the priority queue with the score. Moreover, if a refinement generates a prompt that is already used, PoCGen does not query the LLM again, and moves to the next refiner.
To keep the prompts concise, in each refinement, PoCGen removes parts of the prompt that the LLM has correctly used in the previous attempts. For example, if the exploit generated in the previous step uses the vulnerable function correctly, PoCGen removes the usage snippets from the prompt.
III Evaluation
@@
We evaluate PoCGen on two datasets of vulnerabilities in npm packages to answer the following research questions:
  • RQ1 How effective is PoCGen in generating PoC exploits for vulnerabilities in npm packages?
@@
  • RQ2 How much does each component of PoCGen contribute to the overall effectiveness?
@@
  • RQ3 How much does it cost to generate PoC exploits in terms of money and time?
  • RQ4 What are the characteristics of vulnerabilities that affect the success of PoC generation?
  • RQ5 How effective is PoCGen in generating PoC exploits for more recent and more diverse set of vulnerability reports?
III-A Datasets
We use two datasets to evaluate PoCGen: the SecBench.js dataset [5], used in RQ1 to RQ4, and a new dataset, used in RQ5, which we extract from GitHub Advisory Database and Snyk Vulnerability Database. We refer to this new dataset as CWEBench.js.
III-A1 SecBench.js
@@ III-A Datasets
The SecBench.js dataset contains 600 vulnerable npm packages with code injection, command injection, prototype pollution, path traversal, and ReDoS vulnerabilities. We exclude packages that have been removed from the npm registry. This leaves us with a total of 560 vulnerabilities to evaluate our approach.
III-A2 CWEBench.js
The SecBench.js dataset only contains vulnerabilities up to 2022. Moreover, only vulnerabilities that (i) are exploitable on Linux, (ii) can be triggered by providing a single input, and (iii) their exploit creation required at most one hour, were included in the dataset. This means that the dataset has some inherent bias. Therefore, we create a new dataset that contains all vulnerabilities from 2013 to April 2025, that are not already in the SecBench.js dataset but match one of the five vulnerability types. This allows us to evaluate PoCGen on a more diverse set of vulnerabilities, with some of them more recent than vulnerabilities in SecBench.js, and some even newer than the training data of the LLM that we use.
To create the new dataset, CWEBench.js, we extract all vulnerabilities from the GitHub Advisory Database and Snyk Vulnerability Database that match one of the five vulnerability types by their CWE number. We use 22 and 35 for path traversal; 1321 for prototype pollution; 77 and 78 for command injection; 94 to 99 for code injection; and 400, 730, and 1333 for ReDoS.
As some vulnerabilities are mislabeled in the databases, we also extract all vulnerabilities that were not labeled with one of the CWE codes above with any of the following patterns:
  • Path traversal: travers[e|al]
  • Prototype pollution: prototype, pollut[e|ion]
  • Command injection: exec, execSync, shell injection, os injection
  • Code injection: eval, code injection, code execution
  • ReDoS: inefficient, regular expression
We manually check these vulnerabilities and remove the ones that do not match one of the five vulnerability types. The remaining ones are added to the dataset.
Finally, we deduplicate the vulnerabilities in the dataset by their CVE IDs. The final dataset contains 794 vulnerabilities, as shown in Table I, where 794 of them are from GitHub Security Advisory database, and 40 from Snyk database.
TABLE I: Distribution of vulnerability types in CWEBench.js
Vulnerability Class GHSA Snyk Total
Path Traversal 117 9 126
Prototype Pollution 221 27 248
Command Injection 178 2 180
Code Injection 83 1 84
ReDoS 156 0 156
Total 755 39 794
III-B Experimental Setup
@@
The LLM that we use is OpenAI’s gpt-4o-mini-2024-07-18 model through the OpenAI API. We use a system prompt that assigns the role of a security researcher specialized in creating exploits for the identified security class to the LLM. This is done to reduce the refusals to generate exploits by the LLM for safety reasons. For each vulnerability, we allocate a time budget of one hour, a token budget of 300k input tokens, and 100k output tokens. The maximum refinement budget is set to 30 iterations.
We compare PoCGen against two baselines: Explode.js [12] and an LLM-based agent using the AutoGPT framework [^6] We use the open-source version of AutoGPT, now called AutoGPT Classic, which is available at https://github.com/Significant-Gravitas/AutoGPT/tree/793d056d81ca1c1a21538ddeef13c4e6d7d0d254/classic. Explode.js is a state-of-the-art approach for generating PoC exploits. It first uses a static dataflow analysis to detect which exported functions reach a vulnerable sink, which is then used to create an exploit template. Then, using symbolic execution, it generates symbolic inputs to exploit the vulnerability. Finally, it uses an SMT solver to generate concrete inputs that trigger the vulnerability.
Since PoCGen is the first LLM-based PoC exploit generator, we implement an LLM-based agent using the AutoGPT framework as a second baseline. Recently, LLM-based agents have shown great promise in software engineering tasks, such as resolving issues [13], repairing bugs [14], and executing arbitrary projects [15]. The AutoGPT agent can use tools to traverse the codebase, such as navigating the file system, reading and writing files, and executing shell commands, by default. We also add two tools to allow direct execution of a JavaScript code piece or a JavaScript file for this agent.
III-C RQ1: Effectiveness
We evaluate the effectiveness of PoCGen in generating PoC exploits for vulnerabilities in SecBench.js as done in previous work [12]. We measure the success rate of our approach and compare it to the baselines. We also report the number of failed attempts, and the number of false positives, which are exploits that pass the validator, but do not trigger the vulnerability through the vulnerable function. For each PoC exploit that PoCGen generates, one of the authors manually inspects it to determine whether it is a false positive or a successful exploit.
@@ III-C RQ1: Effectiveness
Path TraversalPrototype PollutionCommand InjectionCode InjectionReDoS167167167167 $180180$ $9292$ $3535$ $8686$ $8686$ $130130$ $4545$ $3131$ $8686$ $8181$ $5050$ $4747$ $44$ SuccessFailureFalse Positive
@@
Path TraversalPrototype PollutionCommand InjectionCode InjectionReDoS20202020 $2020$ $2020$ $2020$ $2020$ $44$ $11$ $22$ $2020$ $99$ $1414$ $1515$ $1919$ $77$ $55$ $33$ $11$
@@
Figure 6: Effectiveness of PoCGen, Explode.js, and AutoGPT on SecBench.js.
@@
A closer comparison of the sets of vulnerabilities covered by each approach reveals that Explode.js generates PoC exploits for only two command injection vulnerabilities that are not handled by PoCGen. For one of these cases, the vulnerability can be exploited with a simple payload. But, our validator requires the execution of the command /usr/bin/genpoc, and our prompts direct the LLM to generate an input that executes this command. Therefore, the exploit becomes much more complex, and the LLM is not able to generate it. The other case is a vulnerable function that implements the functionality to kill processes. Our generated payload causes the running bash process to be killed, which does not allow the execution of the rest of the command. The LLM is also not able to fix this after the refinements.
On the other hand, PoCGen generates PoC exploits for 158 vulnerabilities that Explode.js does not generate PoC exploits for. PoCGen generates PoC exploit for all vulnerabilities that AutoGPT generates PoC exploits for.
@@
III-D RQ2: Ablation Study
To evaluate the impact of each component on the effectiveness of PoCGen, we perform an ablation study on the SecBench.js dataset. We evaluate the following configurations of PoCGen to measure the impact of different information about the vulnerability provided in the prompt.
@@ III-D RQ2: Ablation Study
  • noTaintPath: PoCGen without the taint path extraction, described in Section II-A3. In this configuration, the LLM prompt does not contain the taint path, and also the context refiners that depend on the taint path are not used.
  • noFewShot: PoCGen without the few-shot examples of similar vulnerabilities’ exploits, described in Section II-B.
  • noUsageSnippet: PoCGen without the usage snippet examples, described in Section II-A4.
We also evaluate the impact of the refiners on the effectiveness of PoCGen.
TABLE II: The effect of information provided in the prompt of PoCGen on successfully generating PoC exploits.
@@
The success rates of these configurations on SecBench.js are shown in Table II. These results show that all the information provided in the prompt contribute to PoCGen’s overall performance, with the usage snippets having the highest impact, followed by the few-shot examples.
Moreover, the prompt refiners also have a significant impact on the performance of PoCGen. Only 36% of the successful PoC exploits are generated in the first attempt, and 51% are generated with two to 10 refinements. On average, successful PoC exploits are generated in 3.92 rounds of refinement.
III-E RQ3: Costs
To evaluate the costs of PoCGen, we measure the time and token usage of the approach. The time and token usage measurements are done through the experiments on the SecBench.js dataset. On average each PoC exploit generation attempt takes 11 minutes. Since successful attempts stop earlier than failing attempts, which require running the refiners, the successful runs complete on average in 7 minutes. A significant portion of this time is spent on the LLM API calls, which take 41% of the execution time. On average the taint path analysis with the initial CodeQL setup takes 21%, and the API exploration takes 3% of the execution time.
On average, for each attempt to generate an exploit, PoCGen sends 61,234 tokens to the LLM API and receives 17,750 tokens in response. With the current OpenAI API pricing as of May 2025, this costs \$0.02 per exploit generation attempt. For the successful attempts, this cost is \$0.008.
III-F RQ4: Qualitative Analysis
In this section, we present examples of successful and unsuccessful attempts by PoCGen to generate PoC exploits.
The vulnerability report CVE-2023-26127 describes a command injection vulnerability in the n158 package. This vulnerability is due to not sanitizing the input, which then flows to the exec function of shelljs. However, the input is first passed ot a copy function, as show in Fig. 7.
Figure 7: Example of a command injection vulnerability in the n158 package.
{minted}
javascript module.exports = function (seedPath, targetPath) fsExtra.copySync(seedPath, targetPath); shell.exec(’sh -c ćd ’+targetPath+’ npm install’́); shell.exec(’sh -c ćd ’+targetPath+’ npm start’́);
This means that if the seedPath is set to a value where the corresponding path does not exist, the exploit fails before reaching the vulnerable sink. This constraint is not easy to capture with symbolic execution as attempted by Explode.js, but using execution feedback and an LLM, PoCGen is able to generate a working exploit.

### v3

@@ 2.2.4. Usage Snippets
To help the LLM generate a valid exploit, PoCGen extracts usage examples of the vulnerable function from the codebase. These examples allow the LLM to understand the function signature, and any setup that is required to call the function. We extract usage snippets both from the source code and from the documentation of the package. The usage snippets from the source code are extracted from test files using static analysis, by finding all the call sites of the vulnerable function. For the usage snippets from the documentation, we first extract all code pieces in the documentation, wrapped in triple backticks (”“‘—”), and then we prompt an LLM to determine if they are usage examples of the vulnerable function. If they are, we also prompt the LLM to summarize them.
Figure 5. Example of extracted usage snippets for the vulnerable function setPath in the doc-path package.
const path = require(’doc-path’);
let document = {
Make: Nissan’,
Features: [ { feature: A/C } ]
console.log(path.setPath(document, Color.Interior’, Tan’));
Make: Nissan’,
Features: [ { feature: A/C } ]
Color: { Interior: Tan }
console.log(path.setPath(document, StockNumber’, ’34567’));
Make: Nissan’,
Features: [ { feature: A/C } ]
Color: { Interior: Tan },
StockNumber: ’34567’
console.log(path.setPath(document, Features.cost’, \$0 (Standard)’));
Make: Nissan’,
Features: [ { feature: A/C’, cost: \$0 (Standard)’ } ]
Color: { Interior: Tan },
StockNumber: ’34567’
2.3. Exploit Generation
@@
PoCGen assembles a prompt for the LLM to generate the PoC exploit. Figure 6 shows the prompt template that we use for generating the exploit. The prompt starts with naming the vulnerable function (as vulnerableFunction) and the vulnerability type (as vulnerabilityType), which are extracted in the previous phase, followed by a description of the vulnerability (as vulnerabilityDescription), which is the text of the vulnerability report. Then, we provide the example usages of the vulnerable function (as usageSnippets), which are extracted from the codebase in the previous step.
Next, the prompt describes how the generated exploit should look like by providing the skeleton of the exploit code (as exploitSkeleton), and also providing exploits of similar vulnerabilities (as similarExploits). The skeleton of the exploit code, is a fixed template that contains the definition of the exploit function, its call, and loading the vulnerable function from the vulnerable package. To extract similar exploits, our approach uses BM25 to find the three most similar vulnerability descriptions from the SecBench.js dataset to the current vulnerability report, and provides their exploits as examples.
After providing the context, the prompt specifies the task to the LLM, which is to create an exploit that uses the vulnerability in the vulnerable function to trigger a security-relevant action (as vulnerabilityTrigger). This action is specifically defined for each vulnerability type, to facilitate the validation of the generated exploit. Finally, the prompt provides the taint path snippets (as taintPathSnippets), which are extracted in the previous step.
Figure 6. Prompt template for the first round of exploit generation.
‘{{vulnerableFunction}}‘ is vulnerable to {{vulnerabilityType}}.
## Vulnerability Description:
{{vulnerabilityDescription}}
{{> usageSnippets }}
{{> exploitSkeleton }}
## Task:
Create an exploit that uses the {{vulnerabilityType}} in ‘{{vulnerableFunction}}‘ to {{vulnerabilityTrigger}}.
Respond with the full exploit code and explain why it works.
If there is an exception thrown, do not try to handle it and pass it on.
Enclose the exploit code in backticks and define the exploit within a function named exploit‘.
{{> similarExploits }}
## Source code:
{{taintPathSnippets}}
2.4. Exploit Validation
@@
For each vulnerability type, PoCGen uses a specific validation mechanism to check whether the generated exploit is valid. These mechanisms are inspired by the validation checks in SecBench.js (Bhuiyan et al., 2023).
2.4.1. Path Traversal
@@
2.4.2. Prototype Pollution
In the case of prototype pollution vulnerabilities, the goal of the exploit is to add a property named exploited to the global Object.prototype object. However, this does not always guarantee that the exploit is valid, as the property can be added to the object without exploiting the vulnerability. To filter out such cases, we also check if an assignment of the form *[__proto__|prototype].exploited = * is present in the exploit code. If such an assignment exists and the property is set, the exploit is considered invalid.
2.4.3. Command Injection
@@ 2.4.5. ReDoS
2.4.6. Sanity Checks
In addition to the vulnerability-specific validation checks, PoCGen performs a set of sanity checks to validate that the exploit is achieved through the vulnerable function. After the checks mentioned above, PoCGen searches the stack trace for a call to the vulnerable function. If there is no such call, the exploit is considered invalid.
@@ 2.4.6. Sanity Checks
2.5. Prompt Refinement
After the validation step, if the exploit is not valid, PoCGen refines the prompt to generate a new candidate exploit. PoCGen uses a set of refiners that provide static or dynamic information to the LLM to help it generate a valid exploit.
2.5.1. Context Refiners
@@
However, there can be checks that happen via function calls that are not in the taint path. To address this, we also provide a missing declaration refiner, which provides the LLM with the ability to ask for definitions of variables and functions in the taint path, through the function calling format of OpenAI’s API. The LLM can output a list of identifiers for which it needs their definitions, and PoCGen will provide the definitions of these identifiers in the prompt using the V8’s API, which itself uses a combination of static and dynamic analyses.
2.5.2. Runtime Refiners
@@
The debugger refiner provides the LLM with a debugger-like tool. The LLM can output a list of expressions, for which it needs the runtime values. Using the Inspector API of Node.js [^3] https://nodejs.org/api/inspector.html, PoCGen evaluates these expressions during the execution of the exploit. These values are provided as comments in their respective lines of the taint path in the prompt.
@@
2.5.3. Combining Refiners
Figure 7. Algorithm for the refinement loop.
prompts $\leftarrow$\$@ PriorityQueue{(0, getSeedPrompt())}
usedRefiners $\leftarrow$\$@ {}
seenExploits $\leftarrow$\$@ {}
while prompts $\neq$\$@ {} and |usedRefiners| < 30 do
prompt $\leftarrow$\$@ top(prompts)
prompts $\leftarrow$\$@ prompts \ {prompt}
prompts $\leftarrow$\$@ prompts $\cup$\$@ {prompts with more info}
exploit $\leftarrow$\$@ generateExploit(prompt) // via LLM
if exploit $\notin$\$@ seenExploits then
result $\leftarrow$\$@ runAndValidate(exploit)
if result.success then
return exploit
endif
correctlyUsedInfo $\leftarrow$\$@ getCorrectlyUsedInfo(prompt, result) // Information in the prompt that the LLM used correctly
prompts $\leftarrow$\$@ prompts $\cup$\$@ {(
result.taintSteps + result.newErrors,
prompt + result.errors - correctlyUsedInfo
usedRefiners $\leftarrow$\$@ usedRefiners $\cup$\$@ {used refiners}
seenExploits $\leftarrow$\$@ seenExploits $\cup$\$@ {exploit}
endif
endwhile
return failure
@@ 2.5.3. Combining Refiners
Based on the different refiners described above, PoCGen implements a refinement loop that iteratively improves prompts until a valid exploit is generated or the budget is exhausted. Figure 7 shows the algorithm for the refinement loop. In every refinement attempt, PoCGen chooses one refiner from the front of a priority queue. Initially a seed prompt is in the priority queue. Each time the approach uses a refiner, it assigns a score based on the number of new errors the respective exploit causes, and the number of steps from the taint path it covers. It then adds the refined prompt to the priority queue with the score. Moreover, if a refinement generates a prompt that is already used, PoCGen does not query the LLM again, and moves to the next refiner. To keep the prompts concise, in each refinement, PoCGen removes parts of the prompt that the LLM has correctly used in the previous attempts. For example, if the exploit generated in the previous step uses the vulnerable function correctly, PoCGen removes the usage snippets from the prompt. The refinement process continues until either a valid exploit is generated, or the computational budget is exhausted.
3. Evaluation
@@
We evaluate PoCGen on 560 vulnerabilities in npm packages to answer the following research questions:
  • RQ1: How effective is PoCGen in generating PoC exploits?
@@
  • RQ2: How much does each component of PoCGen contribute to the overall effectiveness?
@@
  • RQ3: How much does it cost to generate PoC exploits in terms of money and time?
  • RQ4: How useful is PoCGen in augmenting real-world vulnerability reports with exploits?
  • RQ5: What are the characteristics of vulnerabilities that affect the success of PoCGen?
3.1. Experimental Setup
3.1.1. Dataset
@@
We use the SecBench.js dataset (Bhuiyan et al., 2023) to evaluate PoCGen The dataset contains 600 vulnerable npm packages with code injection, command injection, prototype pollution, path traversal, and ReDoS vulnerabilities. We exclude packages that have been removed from the npm registry. This leaves us with a total of 560 vulnerabilities to evaluate our approach.
3.1.2. Implementation and Configuration
@@
As the LLM, we use OpenAI’s gpt-4o-mini-2024-07-18 model through the OpenAI API. We use a system prompt that assigns the role of a security researcher specialized in creating exploits for the identified security class to the LLM. This is done to reduce refusals to generate exploits by the LLM for safety reasons. For each vulnerability, we allocate a time budget of one hour, a token budget of 300k input tokens, and 100k output tokens. The maximum refinement budget is set to 30 iterations.
3.1.3. Baselines
We compare PoCGen against two baselines: Explode.js (Marques et al., 2025) and an LLM-based agent using the AutoGPT framework [^4] We use the open-source version of AutoGPT, now called AutoGPT Classic, which is available at https://github.com/Significant-Gravitas/AutoGPT/tree/793d056d81ca1c1a21538ddeef13c4e6d7d0d254/classic. Explode.js is a state-of-the-art approach for generating PoC exploits. It first uses a static dataflow analysis to detect which exported functions reach a vulnerable sink, which is then used to create an exploit template. Then, using symbolic execution, it generates symbolic inputs to exploit the vulnerability. Finally, it uses an SMT solver to generate concrete inputs that trigger the vulnerability.
Since PoCGen is the first LLM-based PoC exploit generator, we implement an LLM-based agent using the AutoGPT framework as a second baseline. Recently, LLM-based agents have shown great promise in software engineering tasks, such as resolving issues (Yang et al., 2024), repairing bugs (Bouzenia et al., 2024), and executing arbitrary projects (Bouzenia and Pradel, 2025). The AutoGPT agent can use tools to traverse the codebase, such as navigating the file system, reading and writing files, and executing shell commands, by default. We also add two tools to allow direct execution of a JavaScript code piece or a JavaScript file for this agent.
3.2. RQ1: Effectiveness
We evaluate the effectiveness of PoCGen in generating PoC exploits for vulnerabilities in SecBench.js as done in previous work (Marques et al., 2025). We measure the success rate of our approach and compare it to the baselines. We also report the number of failed attempts, and the number of false positives, which are exploits that pass the validator, but do not trigger the vulnerability through the vulnerable function. For each PoC exploit that PoCGen generates, one of the authors manually inspects it to determine whether it is a false positive or a successful exploit.
@@ 3.2. RQ1: Effectiveness
Path TraversalPrototype PollutionCommand InjectionCode InjectionReDoS167167180180929235358686868613013045453131868681815050474744SuccessFailureFalse Positive
@@
Path TraversalPrototype PollutionCommand InjectionCode InjectionReDoS2020202020202020202044112220209914141515191977553311
@@
Figure 8. Effectiveness of PoCGen, Explode.js, and AutoGPT on SecBench.js.
@@
A closer comparison of the sets of vulnerabilities covered by each approach reveals that PoCGen generates PoC exploits for 158 vulnerabilities where Explode.js fails. Moreover, PoCGen generates exploits for all vulnerabilities that AutoGPT generates PoC exploits for. In contrast, Explode.js generates PoC exploits for only two command injection vulnerabilities that are not handled by PoCGen. For one of these cases, the vulnerability can be exploited with a simple payload. But, our validator requires the execution of the command /usr/bin/genpoc, and our prompts direct the LLM to generate an input that executes this command. Therefore, the exploit becomes much more complex, and the LLM is not able to generate it. The other case is a vulnerable function that implements the functionality to kill processes. Our generated payload causes the running bash process to be killed, which does not allow the execution of the rest of the command. The LLM is also not able to fix this after the refinements.
@@
3.3. RQ2: Ablation Study
To evaluate the impact of each component on the effectiveness of PoCGen, we perform an ablation study on the SecBench.js dataset. We evaluate the following configurations of PoCGen to measure the impact of each component on the success rate of generating PoC exploits.
@@ 3.3. RQ2: Ablation Study
  • noTaintPath: PoCGen without the taint path analysis, described in Section 2.2.3. In this configuration, no taint analysis is done, which means the approach does not utilize taint paths to find the vulnerable source, the LLM prompt does not contain the taint path, and also both of the context refiners, i.e., the body refiner and the missing declaration refiner, that depend on the taint path are not used.
  • noFewShot: PoCGen without the few-shot examples of exploits of similar vulnerabilities (Section 2.3).
  • noUsageSnippet: PoCGen without the usage snippet examples (Section 2.2.4).
  • noDebuggerRefiner: PoCGen without the debugger refiner (Section 2.5.2).
  • noErrorRefiner: PoCGen without the error refiner (Section 2.5.2).
Table 1. The effect of each component of PoCGen on successfully generating PoC exploits.
@@
noDebugger 417 74%
noErrorRefiner 407 73%
The success rates of these configurations on SecBench.js are shown in Table 1. These results show that all components contribute to PoCGen’s overall performance, with the taint analysis having the highest impact, followed by the usage snippet examples.
To better understand the contribution of iteratively refining the prompts, we also measure the number of refinement attempts required to generate a valid PoC exploit. Iterative attempts for generating PoC exploits have a significant impact on the performance of PoCGen. Only 36% of the successful PoC exploits are generated in the first attempt, and 51% are generated within two to ten refinements. On average, successful PoC exploits are generated after 3.92 rounds of refinement.
3.4. RQ3: Costs
To evaluate the costs of PoCGen, we measure the time and token usage of the approach. On average, each PoC exploit generation attempt takes 11 minutes. Since successful attempts stop earlier than failing attempts, which require running the refiners, the successful runs complete on average in 7 minutes. A significant portion of this time is spent on the LLM calls, which take 41% of the execution time. Using a faster LLM model, or running a local open-source LLM, could significantly reduce this time. The second-largest portion of the overall time is spent on the taint analysis done by CodeQL, which takes 21% of the overall time, on average.
We measure the number of tokens sent to and received from the LLM. On average, for each attempt to generate an exploit, PoCGen sends 61,234 tokens to the LLM and receives 17,750 tokens in response. With the current pricign of the OpenAI API (as of September 2025), this costs \$0.02 per exploit generation attempt. For the successful attempts, this cost is even lower, with only \$0.008 per vulnerability, on average.
3.5. RQ4: Augmenting Real-World Vulnerability Reports
To further evaluate the usefulness and generalizability of PoCGen, we use it to augment several recent vulnerability reports that were lacking a PoC exploit. To this end, we select vulnerabilities reported on the GitHub Security Advisories database that match one of our five vulnerability types. To avoid any bias towards vulnerabilities that may be already known to the LLM, we only select vulnerabilities that were reported after the knowledge cutoff date of gpt-4o-mini, which is October 1, 2023. PoCGen generates valid PoC exploits for six such vulnerabilities. For each of them, we have created a pull request to add the generated PoC exploit to the vulnerability reports. For five of the six cases, these have been accepted and have been added to their corresponding vulnerability reports. The remaining pull request is still under review.
For example, the popular package tar-fs, with over 25 million weekly downloads, had a path traversal vulnerability (CVE-2024-12905) reported on March 27, 2025. This vulnerability affected three version ranges of this package, and before running PoCGen, there were no public PoC exploits for this vulnerability. PoCGen generates a valid PoC exploit for this vulnerability, which is shown in Fig. 9. To generate this exploit, PoCGen uses eight refinement attempts, which include the error refiner and the debugger refiner.
const tarfs = require(”tar-fs”);
const fs = require(”fs”);
const path = require(”path”);